UNH DNP Collaboration Weekly Meeting Minutes

From Nuclear Physics Group Documentation Pages
Jump to: navigation, search

Meeting Minutes

2016-06-02

  • Toby and Ryan waiting on scattering angle from Jie and Pengia; Ellie has asked Jie for weekly updates until it is finalized

Toby

  • Working on dissertation writing
  • Working on creating criteria for a small subset of definitely good runs at at least one setting
  • Then he will attempt to create one good, continuous spectrum and use that to help isolate further issues
  • Planning on presenting this work at next g2p meeting

Ryan

  • Finishing radiative corrections document
  • Talk at JLab was well received
  • Found interesting phenomena where a small phi cut seems to eliminate most of nitrogen and He QE from proton elastic data
  • Will work on angle slices to see how effect changes, whether it can be used to better our packing fraction, and will present at next g2p meeting

Nathaly

  • Accepted into the DNP Summer School in Switzerland!
  • Ellie is working with her, Maurik, Mark McConnell, and Kelly Shaw to handle funding and travel planning before flight prices become expensive
  • Alberto reached out to her about attending the Gordon Conference. She's going to talk with him more about whether GRC can provide some support if she's able to go (might overlap with DNP Summer School)
  • Still at HUGS for the next 3 weeks, Ellie is working with Maurik & Mark to develop a game plan so that she can continue her work on building the fridge shells after she returns

Shujie

  • Currently at HUGS, just moved to JLab for the year to prepare for SRC experiment
  • The SRC group has reached out to her to offer research support, and Ellie is working with Doug and will be joining SRC meetings starting next week to continue to support her

Nick

  • Finished prints for cooling fan mount for 3D printer and has modified the electronics' control code so that the fan runs autonomously as needed for printing
  • Next focus is working on designing and printing a dual-extruder mount to add our new E3D extruder

Ellie

  • Azz. Azz. Azz.
  • Met with Mark and Maurik following news of Patricia's passing
  • Will stay in close contact with them, our goal is to make sure especially that Shujie and Nathaly are supported, and to keep things running so that Karl can take as much time as he needs to focus on his family without worrying about things in the lab
  • Department definitely has funding to cover Shujie and Nathaly through the end of the calendar year, and Ellie is working closely with both of them to make sure all of their needs are being met and their concerns addressed
  • Waiting on hearing back from Mark & Maurik from DOE for long-term plans, will update when there is news
  • Chris Theo has reached out to me about coming back to work for the summer
  • He has work-study funding up to $2500 for the summer
  • Would like to work 30 hr/week for $10/hr, which would cover just over 8 weeks
  • Ellie is working with Maurik on bringing him into the lab to move forward on building circuit boards for the Q Meter power supply, and will be meeting with him next Wednesday
  • Have not heard from Brent, will be following up with him soon to touch base and help him to get started with the NMR

2015-04-29

Toby

  • Tech note has been sent around for comments
  • Beginning to work on asymmetries again, reviving old code to use with updated dilutions

Ryan

  • Radiative cs all set up, working on adding MAID

Nataly

  • Stepper motors arrived, will be installing by next week

Caitlyn

  • Having issues with LabView, will meet with Ellie to address them

Makenzie

  • Tasked with tuning the Q meter to diagnose the recent issue with the modulator output


2015-03-17

Toby

  • Working on dilution factor tech note
  • Has packing fraction through dilution analysis, will work on comparing with Melissa's

Ryan

  • Working on Bosted fits to data, with a focus on the dip region
  • Integrates of dip, DIS, QE, then everything together
  • Good chi^2/dof minimization
  • Looking at changing dip fit to be dependent on Q^2 instead of W to make more global fit parameters
  • Will be presenting results at Wed. g2p meeting

Karl

  • Budget from DOE is finalized
  • Possible trip to JLab to work on their DNP

Ellie

  • Spectator Tagging workshop was good, mEIC folk are very interesting in seeing how the tensor program develops
  • Work on the Azz proposal is on-going, will be moving meetings to MWF schedule
  • Installed magnet safety light and reed switch

2015-03-10

Ryan

  • Working on fitting data with Bosted model
  • Bosted fits dip emperically only!
  • Uses a gaussian so I will try and tune just dip parameters to give best chi-sq
  • Vince said He3 data will have elastic tail already subtracted when he sends it to me

Toby

  • Working on 1.7 GeV dilution (just started running the code)
  • Looked at 1.1 GeV dilution
  • much trickier because we took data on two different ammonia cell lengths
  • and don't have a full set of dilution runs at a lot of momentum settings
  • low Q2 of this setting also straining the model

Caitlyn

  • Having problems doing the wing fit and subtraction in lab view
  • Suggested for her to look at the old PDP for guidance
  • Also suggested to trying doing the fit over a saved data file first and not "live" in lab view

Makenzie

  • Yale card is showing first signs of life, but needs work in understanding it more fully
  • Still working on getting an output from the Yale card
  • Suggested it might be a problem with the gains because the circuit doesn't seem to respond to them

2015-03-03

Ryan

  • Bosted fits work well for Carbon, but not for Vince's nitrogen data
  • Will work ok for one setting, but it has to be retuned for each to the extent that it's not helpful
  • Delta and QE peaks keep shifting and no set of scaling factors applies globally
  • Will be switching to QFS to see if it works better, while waiting on Vince to double-check the nitrogen data
  • Will run next meeting, since Ellie will be out of town

Toby

  • Put event-by-event PID cuts into dilution analysis, which cleaned it up
  • Dilution at 5T, 2.2 GeV transverse looking good
  • Working on tech note summarizing efforts, will be presenting on Wednesday's g2p meeting

Caitlyn

  • Having trouble using old data as an example back in labview
  • Going from labview-->spreadsheet is easy, spreadsheet-->labview is trickier
  • Fridge is open, so she can use live Q curves using the oscillators to test her wing fit subtraction
  • Already signed up for URC

Makenzie

  • Yale card is showing first signs of life, but needs work in understanding it more fully
  • Diode and mod signals getting through, but DC-offset and gain controls not working
  • Issue seems to be in details of grounding, which are in the process of being diagnosed
  • Built a simple housing for it to keep wiring cleaned up
  • Already signed up for URC

Ellie

  • Tufts talk went well, might have convinced an undergraduate to give up biology for nuclear physics
  • Worked heavily on making elastic parameterizations of recent A, B, and t20 data to include in rates code, still in progress
  • Will be working on Spectator Tagging talk and U. of Central Arkansas seminar this week, and will be attending the workshop early next week
  • Found and fixed problem with EG4 vertex angle calculations (phi was using the wrong x/y ratio so that phi didn't change with CLAS sector)
  • Run-by-run analysis of the fix is still in progress

2015-02-24

Ryan

  • Still working on Bosted optimization code.
  • Works well for one setting, developing way to average results over multiple settings so that one fit is applicable over a wide range
  • Working on wisely creating loops for chi^2 optimization

Toby

  • Has been working on dilution analysis for 2.2 GeV, 2.5 T runs
  • This is the setting that Melissa has a packing fraction for. Varying packing fraction by ~5% causes ~30% change in dilution factor, so nailing that is very important
  • 2.2 GeV, 5 T for long. & trans. initial dilution analysis has results
  • Don't match well to the Bosted model yet, so he and Ryan are working together on that to pull out reasonable packing fractions until we have better results from data

Caitlyn

  • Working on Q-curve wing fit, has a code that seems initially to be working
  • Will spend this week doing checks that the code does what it's supposed to

Makenzie

  • Yale card powered and hooked up, but having trouble seeing anything. Will be trouble-shooting the issue throughout the week.
  • Electrically isolated the wire racks under the lambda/2 cable to reduce noise in the Q-meter

Peter

  • Didn't do the deflate-gate test because he had already done it on his own a couple months ago
  • Updated the iTC firmware to hopefully fix the GPIB timeout problem
  • (The firmware update did not fix the GPIB timeout problem. Ellie is in communication with Oxford about it.)

Ellie

  • Big thanks to Ryan, Toby, Peter, & Makenzie for helping to install the 8" piping from the roots pump to the helium fridge (Note: Thick yellow strap has since been added to the green straps at the 1/2-way point)
  • Tested pump line without the fridge, it pulls vacuum well
  • Ran diagnostics of new iTC firmware, confirmed that it did not fix our problem
  • Worked on Hall A gemc simulation for Ay0 paper
  • Reading up on past t20 measurements to determine if any extra equipment is needed to get t20 along with Azz
  • Began updating Azz code to include t20 elastic kinematics, work in progress
  • Heard back from KNF, says that 50 Hz is needed for pump to run as designed
  • Talked with Maurik and representative from Phoenix Phase Converters (who make 1-phase-to-3-phase converters) about what would happen if we used the line we have. Thoughts were that the pump would run faster, probably reducing it's life in the long term. Will be talking with Don Crabb about how their KNF pump is wired.

2015-02-17

Makenzie

  • No lab update, due to the many snow days. Professors are playing catch-up and putting more homework than normal on him, and he's hoping to get back to normal in a week or two.

Toby

  • Continuing work on Dilution. There's an issue around 600 MeV that's still being looked at, as well as why including more corrections moves the dilution away from the model and flattens it out. Work in progress, will update at g2p meeting next week.

Ryan

  • Looking at yield drifts correlated to beam drifts. LHRS and RHRS are drifting in the same direction, which is unexpected. Thinks there is a possible beam scraping far upstream, causing a hot-spot to not be visible. Work is in progress.
  • Redoing and automating his Bosted model code to help Toby with the dilution process

Caitlyn

  • Working on fitting the wings of the Q curve.

Ellie

  • Called Pfieffer and requested a wiring diagram of the pumps, still waiting for it. Called and e-mailed KNF about the 50 Hz pump issue, still waiting for a response. Fridge work on hold until parts come in. Working on EG4-g2p kinematic matching, and including D(e,e'd) t20 calculations to Azz rates. Talk at Tufts got pushed back a week to March 2nd due to snow re-scheduling.